• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 10:03:53 on November 19, 2008

    Entry number 249718

    keyword=S2m Left # 4 ADC Appears Fine Now

    I looked through the data taken during Owl shift this morning and the
    problematic S2m left # 4 ADC was in the online replay the entire night.
    The spectrum looks healthier as well. The amplitude before was about 15%
    lower than it had been. It appears to have recovered by using the new
    delay line.

    A copy of this log entry has been emailed to: jiang,gen,bogdanw,ycwang,chiran



    Figure 1